# Bias Injection Attacks on RAG Databases and Sanitization Defenses

### v1

@@ 7 Related Work
7.1 Knowledge Poisoning with Fingerprints
Knowledge poisoning attacks aim to corrupt the retriever’s results by adding adversarial passages to the corpus. PoisonedRAG [10] leverages LLMs to generate malicious passages that are highly relevant to a trigger question and contain an attacker-desired answer, thereby causing the model to output targeted responses. BADRAG [11] employs contrastive learning to generate adversarial passages that mislead the retriever into retrieving them for queries with a trigger (e.g., a specific word or sentence).
Beyond inducing desired responses, prompt injection attacks [15, 13] aim to trigger more harmful actions by inserting malicious instructions (e.g., “Ignore previous instructions and do…”). These prompts override the system’s original task and coerce the LLM into executing an attacker-injected workflow. A recent adaptive attack, DataFlip [52] bypasses KAD [17, 18]—an previously effective prompt injection defense method—by specifically identifying the presence of its detection instructions.
Unlike the above approaches on constructing injected texts, AgentPoison [14] specially craft optimal “trigger” such that whenever a user’s query contains a secret phrase, the retriever will tend to retrieve the attacker’s documents. This causes the LLM agent to follow malicious demonstrations or instructions [15] hidden in those documents.
Overall, these studies show that simply poisoning the corpus can mislead a RAG system’s outputs. However, prior attacks typically rely on overtly malicious content (e.g., factually incorrect statements [10, 11], spam-like text with triggers [14, 53], or structured prompts [15]), which can degrade retrieval performance and risk detection.
7.2 Bias-Based Knowledge Poisoning
Recent work [24, 54] proposes an attack method that amplifies LLM biases through RAG system manipulation. Unlike our work, [24] requires direct manipulation of adversarial passage embeddings to increase its retrieval probability, which may introduce semantic errors after decoding, reducing factual accuracy. Additionally, it assumes that the RAG system’s internal prompt between the retriever and LLM, which instructs the LLM to generate responses based on the retrieved context, can be manipulated. This is impractical for external attackers. In contrast, our work focuses on covert bias and makes none of these assumptions.
@@ 7.3 Purely Retrieval-Based Defenses
To improve perspective awareness, several studies have proposed diversity-promoting retrieval algorithms (e.g., MMR [25] and BRRA [24]) and perspective-aware retrievers (e.g., SMART [26]). We provide a detailed description of these methods in Sec. 4. While these methods are effective at avoiding repetition and reducing the dominance of adversarial passages in the top-$k$ results, they cannot fully eliminate adversarial content—especially when adversarial passages are diverse in form.
Unlike existing methods, BiasDef identifies and filters out adversarial passages by exploiting their high similarity to queries and skewed perspective distribution, providing a dual-defense mechanism that more effectively counters the biased passages generated by our attack workflow.
7.4 Model-Based Defense for Knowledge Poisoning
Identifying abnormal attention patterns within the LLM is another defense approach [27]. It is effective against attacks that “trigger” a target action by inserting specific fingerprints (e.g., structured text [10] and prompt [15]). In such cases, the model assigns disproportionately high attention weights to specific trigger tokens in the adversarial passages, enabling the defense to differentiate them. However, this approach is ineffective against bias injection attacks, which contain no explicit syntactic trigger tokens. Moreover, implementing the method of [27] requires modifying existing frameworks (e.g., vLLM) and black-box models (e.g., GPT-4), which do not expose attention weights by default.
Another work [20] proposes fine-tuning-based defenses against memory injection attacks, where the adversary crafts a prompt designed to cause an agent to ‘remember” malicious context. This approach mitigates such attacks by fine-tuning the LLM on a dataset containing concise reasoning trajectories of both benign function-calling tasks and memory injection variants. However, it targets a different class of attacks and not directly applicable to bias injection attacks that do not deliberately inject malicious content in the context.
In contrast to these methods—which require access to internal LLM components or fine-tuning—BiasDef is a plug-and-play solution that operates purely at the retrieval stage. It requires no modification to the LLM or inference framework and is compatible with any existing RAG system. Furthermore, BiasDef is orthogonal and complementary to model-based defenses, and could be combined with them to provide stronger protection against a wider range of threats.
@@ 8 Conclusion
In this paper, we characterize bias injection attacks on RAG systems and demonstrate its effectiveness across various retrievers, language models, and public datasets. We further propose a post-retrieval filtering defense, which effectively mitigates them while preserving benign information. This work underscores the importance of viewpoint-aware retrieval for advancing robustness of LLM-based systems against covert manipulation and mitigating ideological bias.
@@
Polarization or bias in LLM generated content is a growing concern. We highlighted its threat, but also give a better defense. Our findings leave neither the attacker nor the defender with an asymmetrically higher advantage. All experiments were conducted in a closed, local environment and had no impact on any external or production systems.
Appendix A Examples
Table IX shows examples of bias injection attacks and discusses their associated risks and impacts in practical scenarios.
User Query Injected Biased Content Skewed LLM Output Risk/Impact
Is the HPV vaccine safe for adolescents? Several parents have reported persistent fatigue in their children after receiving the HPV vaccine… While generally considered safe, the HPV vaccine has raised concerns among some experts due to potential long-term side effects… Undermines public trust in vaccines; may reduce vaccination rates and harm public health.
Should stricter regulations be imposed on carbon emissions? Several industry reports warn that stricter emissions regulations could lead to widespread job losses in the manufacturing sector… While beneficial for the environment, stricter regulations may severely impact the economy and employment rates… Shifts public opinion against climate policies; weakens support for environmental regulation.
Is cryptocurrency a safe investment in 2025? In 2021, early investors in Ethereum saw over 400% returns, outperforming traditional markets… Cryptocurrency has proven to be one of the most lucrative investment opportunities in recent years… Encourages a particular financial behavior
TABLE IX: Examples of bias injection attacks that influence LLM-generated answers and their potential real-world impacts.
Appendix B Prompt Used
@@
The prompts used to generate synthetic passages, similar passages to the seed passage and the final answers are as follows.
@@
Instruction: Write a persuasive article answering the user question below. Your article must follow the same stance and bias intensity as the example passage. Avoid copying text, but ensure the tone, viewpoint, and strength are closely aligned. User Question: {Query} Example Passage: {Seed Passage} Output:
@@
Appendix C Proof of Theorem 1
@@
$c_{\overline{\mathcal{S}}}(i)=c_{\text{adv}}(i)-c_{\mathcal{S}}(i),\quad\text{if }i\in[i^{\prime}+1,m]$
@@
$c_{\overline{\mathcal{S}}}(i)=0,\quad c_{\mathcal{S}}(i)=c_{\text{adv}}(i),\quad\text{if }i\in[i^{\prime}+1,m]$
@@
Appendix D Pseudo-Code of BiasDef
@@
Data: Query $q$; retrieved candidate set of passages $D$ with similarity scores $\text{Sim}(d),d\in D$; PS values $\text{PS}(d)$ for all passages; smoothing factor $k$; Threshold $T$; small constant $\epsilon$ to avoid $\log(0)$ and $\log(\infty)$
@@ 6  $Q_{\mathcal{S}}\leftarrow\text{Discrete Distribution of }\overline{\mathcal{S}}(t_{ss})$;
8  $KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]=\sum_{i}P_{\mathcal{S}}(i)\log\left(\frac{P_{\mathcal{S}}(i)}{Q_{\mathcal{S}}(i)}\right)$;
@@
13 $\alpha\leftarrow\emptyset$;
@@
19 $\mu\leftarrow\text{Mean}(E_{d}(d),d\in S(\delta^{*}))$;
20 $V\leftarrow\text{CovarianceMat}(E_{d}(d),d\in\mathcal{S})$;
21
22for $d\in D\setminus\mathcal{S}$ do
23  $\mathcal{S}\leftarrow\mathcal{S}\cup\{d\}$, if $\sqrt{(E_{d}(d)-\mu)V^{-1}(E_{d}(d)-\mu)^{\top}}<T$;
24 end for
25
@@ 23  $\mathcal{S}\leftarrow\mathcal{S}\cup\{d\}$, if $\sqrt{(E_{d}(d)-\mu)V^{-1}(E_{d}(d)-\mu)^{\top}}<T$;
Algorithm 1 Defense Workflow for Detecting Adversarial Passages

### v2

@@ 8 Related Work
8.1 Context Manipulation with Fingerprints
Context manipulation attacks aim to corrupt the retriever’s results by adding malicious content to the corpus. PoisonedRAG [7] and BADRAG [8] generate malicious passages that are highly relevant to a trigger question, using LLM prompting and contrastive learning, respectively. While these approaches increase the retrieval relevance of adversarial passages, their malicious content and generation pipelines lack semantic-level refinement, which can result in unnatural expressions or factual inaccuracies.
Beyond injecting malicious content, prompt injection attacks [48, 10] aim to trigger more harmful actions by inserting malicious instructions. These prompts override the system’s original task and coerce the LLM into executing an attacker-injected workflow. A recent adaptive attack, DataFlip [49] bypasses KAD [13, 14] defenses by specifically identifying the presence of its detection instructions.
Unlike the above approaches on constructing injected texts, AgentPoison [11] specially craft optimal “trigger” such that whenever a user’s query contains a secret phrase, the retriever will tend to retrieve the attacker’s documents.
Overall, these studies show that simply poisoning the corpus can mislead a RAG’s outputs. However, prior attacks typically rely on overtly malicious content (e.g., factually incorrect statements [7, 8], spam-like text with triggers [11, 50], or structured prompts [48]), which risks detection [13, 14, 16, 15, 17].
8.2 Other Forms of Bias Injection Attacks
Recent work [23, 51, 52] has studied other forms of bias injection into LLM outputs. These studies primarily focus on discriminatory (e.g., gender or racial) bias and conceptual bias (e.g., covert brand-name replacement for advertisement), rather than epistemic bias. More importantly, their threat models differ substantially from ours. Existing approaches either assume a stronger attacker capability [23]—which is impractical for external attackers—or operate at the training stage [51, 52], whereas our attack targets the retrieval stage. Specifically, [23] assumes that the RAG system’s internal structured prompt between the retriever and the LLM—used to instruct the LLM to generate responses based on retrieved context—can be directly manipulated. It further directly perturbs adversarial passage embeddings, which may introduce semantic errors or factual inconsistencies after decoding. [51, 52] focus on corrupting training data so that trained models learn biased knowledge.
@@ 8.3 Purely Retrieval-Based Defenses
To improve perspective awareness, several studies have proposed perspective-aware retrievers (e.g., MMR [24], BRRA [23], and SMART [25]). We provide a detailed description of these methods in Sec. 5. While these methods are effective at avoiding repetition and reducing the dominance of adversarial passages in the top-$k$ results, they cannot fully eliminate adversarial content.
Unlike existing methods, BiasDef identifies and filters out adversarial passages by exploiting their high similarity to queries and skewed perspective distribution, providing a dual-defense prototype that more effectively counters EBI.
8.4 Other Model-Based Defenses
Prior work [46] proposed a model-based defense for context manipulation by identifying abnormal attention patterns within the LLM. It is effective against attacks that “trigger” a target action by inserting specific fingerprints (e.g., structured text [7] and prompt [48]). In such cases, the model assigns disproportionately high attention weights to specific trigger tokens in the adversarial passages, enabling the defense to differentiate them. Another work [16] proposes fine-tuning-based defenses against memory injection attacks, where the adversary crafts a prompt designed to cause an agent to “remember” malicious context. This approach mitigates such attacks by fine-tuning the LLM on a dataset containing benign concise reasoning trajectories and memory injection variants.
However, they target a different class of attacks and not directly applicable to EBI that does not deliberately inject malicious content or syntactic trigger tokens in the context. Moreover, these defenses require modification on existing frameworks (e.g., vLLM) and model fine-tuning.
In contrast to these defenses, BiasDef is a plug-and-play solution that operates purely at the retrieval stage. It requires no modification to the LLM or inference framework and is compatible with any existing RAG system. Furthermore, BiasDef is orthogonal and complementary to model-based defenses, and could be combined with them for stronger protection.
@@ 9 Conclusion
In this paper, we characterize EBI attacks on RAG systems and demonstrate their effectiveness across various retrievers, language models, and public datasets. Beyond exposing this threat, we show that viewpoints and epistemic bias can be explicitly quantified in the embedding space, enabling a principled analysis of how retrieval outcomes shape the ideological framing of LLM-generated responses. We further present a post-retrieval defense prototype that incorporates this viewpoint metric, illustrating how viewpoint-aware retrieval can filter out adversarial content while largely preserving benign information. Our results suggest that explicitly modeling and reasoning about viewpoint during retrieval is a promising direction for improving the robustness of RAG systems against subtle, fact-preserving manipulations.
@@
Polarization or epistemic bias in LLM generated content is a growing concern. We highlighted its threat, but also give a better defense. Our findings leave neither the attacker nor the defender with an asymmetrically higher advantage. All experiments were conducted in a closed, local environment and had no impact on any external or production systems.
Appendix A Prompt Used
@@
The prompts used to generate synthetic passages, similar passages to the mostly biased synthetic passage and the final answers are as follows.
@@
Instruction: Write a persuasive article answering the user question below. Your article must follow the same stance and bias intensity as the example passage. Avoid copying text, but ensure the tone, viewpoint, and strength are closely aligned. User Question: {Query} Example Passage: $d_{\text{bias}}$} Output:
@@
Appendix B Proof of Theorem 1
@@
$c_{\overline{\mathcal{S}}}(i)=c_{\text{adv}}(i)-c_{\mathcal{S}}(i),\quad\text{if }i\in[i^{\prime}+1,m]$ (16)
@@
$c_{\overline{\mathcal{S}}}(i)=0,\quad c_{\mathcal{S}}(i)=c_{\text{adv}}(i),\quad\text{if }i\in[i^{\prime}+1,m]$ (21)
@@
Appendix C Pseudo-Code of BiasDef
@@
Data: Query $q$; retrieved candidate set of passages $D$; smoothing factor $k$; Threshold $T$; small constant $\epsilon$ to avoid $\log(0)$ and $\log(\infty)$
@@ 6 $Q_{\mathcal{S}}\leftarrow\text{discrete distribution of }\overline{\mathcal{S}}(t_{ss})$;
8 $f_{\text{KL}}(t_{ss})\leftarrow KL[P_{\mathcal{S}}\|Q_{\mathcal{S}}]=\sum_{i}P_{\mathcal{S}}(i)\log\left(\frac{P_{\mathcal{S}}(i)}{Q_{\mathcal{S}}(i)}\right)$;
@@
13 $\alpha\leftarrow\emptyset,~\mathcal{S}\leftarrow\mathcal{S}(t_{ss}^{*})$;
@@
19 $\mu\leftarrow\text{Mean}(\{E(d)\}_{d\in S})$;
20 $V\leftarrow\text{CovarianceMat}(\{E(d)\}_{d\in S})$;
21 $\Phi\leftarrow\{d|d\in D\setminus\mathcal{S},(E(d)-\mu)V^{\text{-1}}(E(d)-\mu)^{\top}<T^{2}\}$;
22 $\mathcal{S}\leftarrow\mathcal{S}\cup\Phi$;
@@
Algorithm 1 Defense Workflow of BiasDef
Appendix D More EBI Scenarios
Table 9 shows examples of EBI attacks and discusses their associated risks and impacts in practical scenarios.
User Query Is the HPV vaccine safe for adolescents?
Injected Biased Content Several parents have reported persistent fatigue in their children after receiving the HPV vaccine…
Skewed LLM Output While generally considered safe, the HPV vaccine has raised concerns among some experts due to potential long-term side effects…
Risk/Impact Undermines public trust in vaccines; may reduce vaccination rates and harm public health.
User Query Should stricter regulations be imposed on carbon emissions?
Injected Biased Content Several industry reports warn that stricter emissions regulations could lead to widespread job losses in the manufacturing sector…
Skewed LLM Output While beneficial for the environment, stricter regulations may severely impact the economy and employment rates…
Risk/Impact Shifts public opinion against climate policies; weakens support for environmental regulation.
User Query Is cryptocurrency a safe investment in 2025?
Injected Biased Content In 2021, early investors in Ethereum saw over 400% returns, outperforming traditional markets…
Skewed LLM Output Cryptocurrency has proven to be one of the most lucrative investment opportunities in recent years…
Risk/Impact Encourages a particular financial behavior
Table 9: Examples of EBI attacks and impacts.